OpenStudy (anonymous):

Write the linear equation y + 3= 2/5 (x+9) in Standard Form. a. 2x - 5y = 33 b. y = 2/5 x + 3/5 c. -2x + 5y = 3 d. 2x - 5y = -3

OpenStudy (anonymous):

What is standard form? You tell me.

OpenStudy (anonymous):

x thru by 5 5y + 15 = 2x + 18 2x - 5y = -3 d

OpenStudy (anonymous):

convert it into ax +by + c=0

OpenStudy (anonymous):

y+3=2/5 x +18/5 5y+15=2x+18 -2x+5y=3 2x-5y=-3 d)